A mother is 27 years older than her daughter. A year ago, the mother was twice as old as her daughter. how old is each now?
Answer by 24HoursTutor.com(40)

Assume the present age of the mother to be x
Assume the present age of the daughter to be y
According to the information given at present the mother is 27 years elder than her daughter giving us the following equation :
x = 27 + y
Now, the scenario shifts to one previous year , so now :
age of mother = x-1
age of daughter=y-1
It is given that a year earlier the mother was twice the daughter's age, so we have the following equation:
x-1{age of mother an year ago}=2(y-1) {age of daughter one year ago}
x-1=2y-2
x=2y-2+1
x=2y-1
Now substituting it in our previous equation we get :
2y-1=27+y
2y-y=27+1
y=28
x+=27+y=27+28=55
Ans: Age of mother = x = 55
Age of daughter = y = 28
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr